OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (anonymous):

y-4=-3(x-1)

iYuko (iyuko):

\[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) m being slope

iYuko (iyuko):

\[y-4=-3(x-1\]